Based on a smartphone​ survey, assume that 40​% of adults with smartphones use them in theaters. In a separate survey of 227 adults with​ smartphones, it is found that 74 use them in theaters.a. If the 40% rate is​ correct, find the probability of getting 74 or fewer smartphone owners who use them in theaters.b. Is the result of 74 significantly​ low?

a. If the 40 rate is​ correct, the probability of getting 74 or fewer smartphone owners who use them in theaters is nothing. ​(Round to four decimal places as​ needed.)
Answer #1

a) Mean = np = 227*0.40 = 90.8

Standard deviation = = 7.381

Hence,

P(74 or fewer)

= P(X < 74.5)

= P(z < (74.5 - 90.8)/7.381)

= P(z < -2.21)

= 0.0136